OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

i have no idea how to do this, but the good news is only one more after this one

OpenStudy (anonymous):

this is my last one, just finished the other

OpenStudy (imstuck):

I really do not understand this at all! I have never seen anything like it in my life! Let me research this a bit. Meanwhile, can you help me with something? I see that you posted two things to me in a row. How do you do that? When I post to someone, even after I click Post, the words are still in this box and a "live preview of your reply" shows and neither one of them goes away. It stays in my box and does not free it up for more typing. How do I delete it after I click post? answer me this while I look up your solution please!

OpenStudy (yoyo123):

This is pretty simple is all you looking for an answer to x and y?

OpenStudy (imstuck):

I thought that, too. Are you just solving for x and y?

OpenStudy (yoyo123):

They have me lost in all that stupid wording english is a weak point of mine. Math is by far my best subject

OpenStudy (anonymous):

no idea

OpenStudy (yoyo123):

well if you want the answer

OpenStudy (yoyo123):

pick one of the systems doesn't matter and solve for x or y

OpenStudy (anonymous):

ok thanks

OpenStudy (yoyo123):

then plug that into the other system

OpenStudy (imstuck):

Does that make sense to you? We can help if you don't understand.

OpenStudy (yoyo123):

I would help you out on this but i can already tell by looking at it there are going to fractions all over the place. and i really dont want to do anything atm

OpenStudy (anonymous):

i cant figure it out -.-

OpenStudy (yoyo123):

-9x + 5y = 16 +9x +9x 5y=9x+16 divide by 5 on both sides y=(9/5)x+(16/5)

OpenStudy (anonymous):

they're both negative 4

OpenStudy (yoyo123):

so y equals that now put it inot the other equation

OpenStudy (anonymous):

at least thats what i got

OpenStudy (yoyo123):

i wish i could show you on paper its a lot easier on paper

OpenStudy (anonymous):

yea its -4

OpenStudy (anonymous):

thanks guys